### v2

@@ 4.4 Discussion
Different ODI SR backbones. In Tab. 6, we compare the performance with different backbones in the ODI SR task. i.e., SwinIR, Omni-SR, and EDSR-baseline. However, we find that by employing deeper SR backbones, the performance of depth estimation drops. For example, with SwinIR as backbone, the Abs Rel metric drops 0.0088 compared with using EDSR-baseline. We think the reason is that the deeper SR backbones are not easy to train. Specifically, the Stanford2D3D dataset only contains 1000 training samples, which is not friendly for attention-based backbones. In addition, the deeper attention-based backbones need more time to converge. Instead, our employed depth estimation backbone, e.g., OmniDepth, would converge in advance in the training process.
Generalization ability. To verify the generalization ability, we further test on 3D60 dataset with pre-trained depth estimation models, i.e., OmniDepth and UniFuse-Fusion, from Stanford2D3D and Matterport3D datasets in Tab. 7. With UniFuse-Fusion as backbone, our weakly-supervised method outperforms the corresponding fully-supervised method, e.g., 3.7% gain and 8.7% gain in MAE metric on Stanford2D3D dataset and Matterport3D dataset, respectively. With OmniDepth as backbone, our method also shows significant improvements, e.g., 10.40% and 0.66% in $\delta_{3}$ metric on Stanford2D3D and Matterport3D datasets, respectively.
@@ 5 Conclusion
In this paper, to the best of our knowledge, we proposed the first work to estimate the HR omnidirectional depth map directly from an LR ODI, when no HR depth GT map is available. We found that the ODIs and their corresponding depth maps share the common scene structural knowledge. Therefore, we introduced an ODI SR task to boost the performance of the HR depth estimation task. Especially, the textural details in the ODI SR task can interfere with the smoothness of depth predictions. In this case, we introduced uncertainty estimation to the ODI SR specially for extracting scene structural knowledge. Upon this, we proposed the CIIF and FD loss in the SSKT module. The CIIF captures the correlations among neighboring pixels and shares the scene structural knowledge between the two tasks, while the FD loss provides extra structural regularization. Experimental results demonstrated that our weakly-supervised method is 1) efficient without extra inference cost, 2) flexible as a plug-and-play approach, and 3) effective to achieve comparable results with the fully-supervised methods.
Limitation and future work: In this paper, the feature up-sampling process in the HR depth estimation task is conducted with integral factors. In practice, predicting the depth maps with arbitrary resolution is required for practical usage. In the future, we will continue to explore how to predict HR omnidirectional depth map with arbitrary resolution from an LR ODI input, using LR depth maps for supervision.

### v3

@@ 4.4 Discussion
Different ODI SR backbones. In Tab. 7, we compare the performance with different backbones in the ODI SR task. i.e., SwinIR, Omni-SR, and EDSR-baseline. However, we find that by employing deeper SR backbones, the performance of depth estimation drops. We think the reason is that the deeper SR backbones are not easy to train. Specifically, the Stanford2D3D dataset only contains 1000 training samples, which is not friendly for attention-based backbones.
Generalization ability. To verify it, we further test on 3D60 dataset with pre-trained depth estimation models, i.e., OmniDepth and UniFuse-Fusion, from Stanford2D3D and Matterport3D datasets in Tab. 8. With UniFuse-Fusion as the backbone, our weakly-supervised method outperforms the corresponding fully-supervised method, e.g., 3.7% gain and 8.7% gain in MAE metric on Stanford2D3D dataset and Matterport3D dataset, respectively. With OmniDepth as the backbone, our method also shows significant improvements, e.g., 10.40% and 0.66% in $\delta_{3}$ metric on Stanford2D3D and Matterport3D datasets, respectively.
@@ 5 Conclusion
In this paper, to the best of our knowledge, we proposed the first work to estimate the HR omnidirectional depth map directly from an LR ODI, when no HR depth GT map is available. We found that the ODIs and paired depth maps share the common scene structural knowledge. Therefore, we introduced an ODI SR task to boost the performance of the HR depth estimation task. Especially, the textural details in the ODI SR task can interfere with the smoothness of depth predictions. In this case, we introduced uncertainty estimation to the ODI SR specially for extracting scene structural knowledge. Upon this, we proposed the CIIF and FD loss in the SSKT module. The CIIF captures the correlations among neighboring pixels and shares the scene structural knowledge between the two tasks, while the FD loss provides extra structural regularization. Experimental results demonstrated that our weakly-supervised method is 1) efficient without extra inference cost, 2) flexible as a plug-and-play approach, and 3) effective in achieving comparable results with the fully-supervised methods.
Limitation and future work: In this paper, we achieve the high-resolution $360^{\circ}$ depth estimation task via weakly-supervised learning. In this way, we alleviate the difficulty of acquiring high-resolution $360^{\circ}$ depth ground truth. However, acquiring a $360^{\circ}$ depth map is inherently difficult due to stitching and alignment processes. Recently, the depth foundation models of perspective depth estimation have risen by exploiting large-scale labeled and unlabeled perspective images [47]. In the future, we plan to excavate knowledge from such rich perspective images and depth maps, which is an appealing direction to inspire the $360^{\circ}$ depth estimation.
.1 Cylindrical Implicit Interpolation Function
.1.1 Implicit Image Function
In general, the image interpolation problem can be defined as follows:
$I_{hr}(x_{q})=\sum_{i\in N_{q}}w_{q,i}v_{q,i},$ (12)
as shown in Fig. 9(a), where $x_{q}$ is the coordinate of query pixel $q$ in the HR image domain, $x_{i}$ is the coordinate of corner pixel $i\in\{00,01,10,11\}$ in the LR image domain and $N_{q}$ is the set of neighbor pixels for $q$, $w_{q,i}$ is the interpolation weight between $q$ and $i$, and $v_{q,i}$ is the interpolation value. $I_{hr}(x_{q})$ is the pixel value of $q$. In traditional image interpolation, the pixel value $I_{lr}(x_{i})$ is commonly used as the interpolation value $v_{q,i}$. Due to the nature of perspective images, $N_{q}$ is usually chosen as the four nearest corner pixels of $q$ in the LR domain (as illustrated in Fig. 9(a), $i\in\{00,01,10,11\}$), and the interpolation weight $w_{q,i}$ is often chosen as the partial area diagonally opposite the corner pixel $i$ (e.g., in Fig. 9(a), interpolation weight $w_{q,00}=S_{00}$), such as Bilinear interpolation.
In implicit image function [48, 32], the image coordinates are normalized into (--1, 1) to make it possible to share the coordinate in the LR image domain and the HR image domain. In this way, [48, 32] can achieve arbitrary up-sampling factors by querying the pixels in the HR image domain according to corresponding coordinates. For example in Fig. 9(a), we can follow the coordinate $(-0.5,-0.5)$ to query the pixel $x_{00}$ in the LR image domain and follow the coordinate $(-0.125,-0.125)$ to query the pixel $x_{q}$ in the HR image domain.
Figure 8: The illustration of cylinder angle coordinate system for ERP images.
Figure 9: (a) General image interpolation process. (b) LIIF [48]. (c) JIIF [32].
Figure 10: The illustration of the up-sampling process from LR feature to HR feature with our CIIF.
Furthermore, as shown in Fig. 9(b), in implicit image functions, instead of directly using the pixel value (e.g., RGB value and depth value), the interpolation value $v_{q,i}$ is often generated through a multi-layer perceptron (MLP) $g_{\theta}$ which takes the latent code $z_{i}$ from an image feature extractor and the relative coordinate $x_{q}-x_{i}$ as input, formulated as:
$v_{q,i}=g_{\theta}(z_{i},x_{q}-x_{i}).$ (13)
In JIIF [32], as shown in Fig. 9(c), instead of utilizing the partial area, the interpolation weight $w_{q,i}$ is predicted together with the interpolation value $v_{q,i}$ through an MLP, which views the interpolation weight and the interpolation value as the edge and the node in a graph model, and can be written as:
$w_{q,i},v_{q,i}=g_{\theta}(z_{i},x_{q}-x_{i}).$ (14)
.1.2 Cylindrical Representation
The aforementioned implicit image functions are designed for perspective images based on Cartesian coordinate system, which are not applicable for omnidirectional images. This is because omnidirectional images are from a sphere, and the most popular equirectangular projection (ERP) is unfold from the sphere based on Cylindrical coordinate system. Using Cartesian coordinate system makes the left side and right side of an ERP image not contiguous. To better represent the spatial correlation of pixels in ERP images, we reframe ERP images with cylindrical representation, as shown in Fig. 8. Note that the range of the cylindrical angle coordinate system ($\alpha,\beta$) is:
$\begin{array}[]{c}\alpha\in(-\arctan(\frac{\pi h}{w}),\ \arctan(\frac{\pi h}{w})),\\ \\ \beta\in(-\pi,\ \pi).\end{array}$ (15)
In this case, our CIIF can be formulated as $g_{\theta}(z_{i},(\alpha_{i}-\alpha_{q},\beta_{i}-\beta_{q}))$. As for the output content, we will discuss it in Sec. .1.3 according to our specific framework and demand.
.1.3 Sharing Strategy
In the above discussion, we have reframed CIIF in cylindrical representation. The aim of CIIF is to transfer structural knowledge from the ODI SR task to the HR depth estimation task. The ODIs and their corresponding depth maps are two different representations for the same scene, which implys that both modals contain the common structural knowledge. Therefore, we directly choose to share parameters of CIIF between the ODI SR task and HR depth estimation task. Note that the parameters of CIIF are all included in an MLP.
As for the output form of CIIF, we find that the predicted interpolation value $v_{q,i}$ in Eq. 13 and Eq. 14 can not be shared between different modals. Instead, the relative correlation among neighboring pixels can effectively represent the common structural knowledge of the same scene, depicted as the interpolation weight $w_{q,i}$. In this case, we only output the interpolation weights $w_{q,i}$ from an MLP, formulated as follows:
$w_{q,i}=g_{\theta}(z_{i},(\alpha_{i}-\alpha_{q},\beta_{i}-\beta_{q})).$ (16)
.1.4 Feature Up-sampling Process
As shown in Fig. 10, given an LR feature map $F_{lr}\in h\times w\times c$, we aim to obtain an HR feature map $F_{hr}\in sh\times sw\times c$, where $s$ is the up-sampling factor. An MLP with hidden dimension $[256,128,64]$ is employed to output interpolation weights $w_{q,i}\in h\times w\times(4*s^{2})$. Then, interpolation weights $w_{q,i}$ is reshaped to $sh\times sw\times 4$. We add in the cylindrical relative angle coordinate with the form $\rm{cos}(\alpha_{i}-\alpha_{q})*\rm{cos}(\beta_{i}-\beta_{q})$. We normalize $w_{q,i}$ in the third dimension by $w_{q,i}=\frac{w_{q,i}}{\sum_{i\in 4}w_{q,i}}$, and thus each query pixel $q$ in the HR domain has four normalized interpolation weights associated with four corner pixels. Then, $F_{hr}$ is obtained with $w_{q,i}$ and $F_{lr}$ based on Eq. 12, where $F_{lr}$ is the interpolation values.
Table 9: The number of samples in training set, validation set, and testing set in three datasets.
Dataset
Stanford2D3D [25] Matterport3D [26] 3D60 [9]
Training 1000 7829
Validation 40 947
Testing 373 2014 1298
.2 FD Loss
We visualize the HR ODI, the HR RGB feature from the ODI SR task, the HR transformed feature and the HR feature from the HR depth estimation task in Fig. 11, respectively. We can find that before transformation, the HR RGB feature from the ODI SR task contains much texture information, which interferes with the smoothness of the HR depth estimation. Instead, the HR transformed feature is more similar with the HR feature from the HR depth estimation task, which is more appropriate for matching the similarity between different tasks.
Figure 11: Visualization of the features employed in the FD loss. (a) HR ODI. (b) HR RGB feature from the ODI SR task. (c) HR transformed feature. (d) HR feature from the HR depth estimation task.
.3 Experiment
.3.1 Dataset and Metrics
Dataset. We follow the split files of UniFuse [^3] https://github.com/alibaba/UniFuse-Unidirectional-Fusion/tree/main/UniFuse/datasets to split the three datasets respectively, as depicted in Tab. 9. As the 3D60 dataset is only used to verify the generalization ability, the splits of its training and validation subsets are omitted.
Metrics. We evaluate our method with standard metrics introduced in [5]. We only calculate the observed pixels $p\in\mathrm{obs}$, which are valid in the HR depth GT maps $D^{\mathrm{GT}}_{\mathrm{HR}}$. Notably, during training, our weakly-supervised method only utilizes LR depth maps $D^{\mathrm{GT}}_{\mathrm{LR}}$ as the supervision, while HR depth GT maps $D^{\mathrm{GT}}_{\mathrm{HR}}$ is only utilized for evaluation. Given $D^{\mathrm{GT}}_{\mathrm{HR}}$ and HR depth prediction $D_{\mathrm{HR}}$, the metrics include:
  • Mean Absolute Error (MAE) measures the absolute difference between the depth prediction and ground truth. It is to show how close between the prediction and ground truth.
    $\frac{1}{\mathrm{N}_{\mathrm{obs}}}\sum_{p\in\mathrm{obs}}||D_{\mathrm{HR}}(p)-D^{\mathrm{GT}}_{\mathrm{HR}}(p)||.$ (17)
  • Absolute Relative Error (Abs Rel) measures the absolute difference between the depth prediction and ground truth, scaled by the ground truth. It is to express how large the error is compared with the ground truth.
    $\frac{1}{\mathrm{N}_{\mathrm{obs}}}\sum_{p\in\mathrm{obs}}\frac{||D_{\mathrm{HR}}(p)-D^{\mathrm{GT}}_{\mathrm{HR}}(p)||}{D^{\mathrm{GT}}_{\mathrm{HR}}(p)}.$ (18)
  • Root Mean Square Error in log space (RMSE log) measures the square root of the squared difference. The squared difference can emphasize the scenes with large depth variances. In addition, utilizing the log form can emphasize scale differences instead of absolute distance differences.
    $\sqrt{\frac{1}{\mathrm{N}_{\mathrm{obs}}}\sum_{p\in\mathrm{obs}}||\log_{10}D_{\mathrm{HR}}(p)-\log_{10}D^{\mathrm{GT}}_{\mathrm{HR}}(p)||^{2}}.$ (19)
  • $\delta_{i}$, the fraction of pixels where the relative error between the depth prediction $D_{\mathrm{HR}}$ and depth GT map $D^{\mathrm{GT}}_{\mathrm{HR}}$ is less than the threshold $1.25^{i}$, $i\in\{1,2,3\}$. It is to evaluate the accuracy ratios rather than absolute or squared differences.
    $max\{\frac{D_{\mathrm{HR}}(p)}{D^{\mathrm{GT}}_{\mathrm{HR}}(p)},\frac{D^{\mathrm{GT}}_{\mathrm{HR}}(p)}{D_{\mathrm{HR}}(p)}\}<i,$ (20)
where $\mathrm{N}_{\mathrm{obs}}$ represents the total number of the observed pixels in $D^{\mathrm{GT}}_{\mathrm{HR}}$.
Table 10: Ablation studies on Berhu loss.
     Abs Rel $\downarrow$      $\delta_{3}\uparrow$
     L1 loss      0.1609      97.18
     Berhu loss [40]      0.1585      97.37
Table 11: Ablation studies on the effect of spherical convolutions.
EquiConv MAE $\downarrow$ Abs Rel $\downarrow$ $\delta_{1}\uparrow$
0.3144 0.1589 79.16
0.3135 0.1585 78.06
Table 12: Investigation of applying CIIF to the fully-supervised methods.
MAE $\downarrow$ Abs Rel $\downarrow$ $\delta_{1}\uparrow$
Conv 0.3124 0.1582 80.16
CIIF 0.3049 0.1577 79.66
Table 13: Ablation studies on FD loss about different calculations of feature similarity and whether applying transform.
Transform Abs Rel $\downarrow$ $\delta_{3}\uparrow$
Feature affinity [37] 0.1645 96.93
0.1630 97.42
Cosine similarity 0.1723 97.34
0.1585 97.37
Table 14: Discussion about the effect of employing SwinIR, which is a deeper SR backbone.
Setting Abs Rel $\downarrow$ RMSE log $\downarrow$ $\delta_{3}\uparrow$
3 0.3414 0.2630 75.25
5 0.4717 0.2653 77.73
Table 15: Comparison of computational costs and time consuming for feature up-sampling between Bicubic interpolation and our CIIF.
        CIIF         Bicubic
        GFLOPs         0.65         1.21
        Time (ms)         26.5         0.5
Table 16: Comparison of computational costs among different input sizes and output sizes.
Depth Backbone Input Size Output Size GFLOPs
OmniDepth [9] $128\times 256$ $512\times 1024$ 21
$512\times 1024$ 317
$256\times 512$ $1024\times 2048$ 86
$1024\times 2048$ 1267
UniFuse-Fusion [5] $128\times 256$ $512\times 1024$ 6
$512\times 1024$ 63
$256\times 512$ $1024\times 2048$ 22
$1024\times 2048$ 250
Figure 12: Illustration of the “noise problem”. (a) The predicted HR depth map in Setting 1; (b) the predicted HR depth map in Setting 3; (c) the LR input in Setting 1; (d) the super-resolved ODI in Setting 3.
.3.2 Ablation Studies
Berhu loss. In Tab. 10, it can be seen that employing Berhu loss [40] has an improvement in both metrics, such as 0.0024 in Abs Rel metric, which can benefit the depth estimation performance.
Spherical convolutions. Recently, some efforts have been made to adjust the shape [9] and sampling location [11] of convolution filters to address the distortion problem in ERP type ODI. To verify if the spherical convolution benefits our framework, we choose EquiConvs, which is characterized by their irregular convolutional kernels to adapt to the varying distortions across different latitudes. However, as introduced in BiFuse [7], the effectiveness of EquiConvs is uncertain, especially in deeper networks. As shown in Table 11, the results show that the influence caused by EquiConvs is minor (0.0009 drop in MAE, and 1.10% improvement in $\delta_{1}$). Therefore, it is reasonable that our design in knowledge transfer is more important than applying the EquiConvs.
FD loss. We further investigate the choice of different calculations of feature similarity in Tab. 13. Without transform, the feature affinity loss [37] even performs worse than cosine-similarity. After adding the transform process, the feature affinity loss improves in both metrics, such as 0.49% in $\delta_{3}$ metric. The results show that the transform process is more important than the selection of the calculation of feature similarity.
.3.3 Discussion
Adding CIIF to current methods. In the main paper, the fully-supervised methods are learnable up-sampling convolutional layers, such as combining Bicubic up-sampling and convolution layers, which is used in UniFuse. In Tab. 12, we try to replace the learnable up-sampling layers with our CIIF. We find that the performance of the fully-supervised can be improved significantly, such as 0.0075 in MAE metric. It demonstrates the effectiveness of the CIIF.
Different ODI SR backbones. In Setting 3 and 5, the choice of deeper SR backbones might provide more accurate SR ODI for the followed depth estimation network. However, In Tab. 14, we find that the depth estimation results in Setting 3 and 5 are still bad. To better illustrate the noise problem, we provide the visualization results in Fig. 12. We observe that the predicted HR depth map from the LR ODI in Setting 1 is obviously better than the predicted result from the SR ODI in Setting 3. In contrast, our framework do not utilize the SR image, but utilize the SR features for knowledge distillation.
Computational costs and inference time. We further report the computational costs with different input sizes and output sizes in Tab. 16. With OmniDepth as backbone, to achieve the goal of HR depth estimation with $1024\times 2048$, our method needs 86 GFLOPs. However, with the input of the original resolution, the computational costs increase to 1267 GFLOPs, which is about 15 times of our method. It demonstrates that our method relieves the computational burden with LR ODI as input during inference.
As shown in Fig. 13, we calculate computational costs with ×$\times 8$ up-sampling factor, EDSR-baseline [41] as the feature extractor in the ODI SR task, and UniFuse-Fusion [5] as the feature extractor in the HR depth estimation task. During training, the feature extractor in the ODI SR task has 9.829 GFLOPs, and the feature extractor in the HR depth estimation task has 0.977 GFLOPs. CIIF which serves for up-sampling feature maps consumes 0.654 GFLOPs. The decoder of each modal is also reported. Once training is done, the ODI SR task can be freely removed, introducing no extra inference cost. Therefore, the overall inference cost is only 1.706 GFLOPs.
.3.4 More visual results.
In Fig. 13, we provide more visual results on Matterport3D dataset. In addition, in Fig. 15, we provide visual comparisons about ablation studies.
Figure 13: Overview of computational costs of each module.
Figure 15: Visual comparison of the impacts of uncertainty estimation in the ODI SR task and FD loss on the overall framework.